1. ANDROGEL 4.12% w/w is a testosterone gel applied topically in males for endogenous testosterone deficiency. For a starting dose of 22.3 mg testosterone, calculate the quantity, in grams, of gel administered.

2. A morphine phosphate extended-release liposome injection contains morphine sulfate 24.5 mg/5mL of injection. Calculate the percent strength of morphine sulfate in the injection.
Answer #1

ANSWER FOR Q.1

At first we have to know that Gel is a colloidal solution which has liquid dispersed phase and solid dispersion medium.

for ANDROGEL we can assume it as a 4.12% w/w colloidal solution.

in which testosterone is dispersed phase ( as it is a hormone and hormone can't be solid)

From the data it is seen that

4.12 gm testosterone present in 100 gm ANDROGEL

0.0223 gm testosterone present in = (100/4.12) * .0223 gm ANDROGEL ( AS 22.3 mg = .0223 gm )

= 0.541 gm ANDROGEL

So 0.541 gm quantity of Gel administrated

ANSWER FOR Q.2

MORPHINE SULPHATE injection contains MORPHINE SULPHATE PENTAHYDRATE

Molecular Wt. of MORPHINE SULPHATE PENTAHYDRATE is 758.833 gm/ mole

that means at NTP 22400 mL of MSP CONTAINS 758.833 gm

758833 mg MSP has volume 22400 mL (MSP = MORPHINE SULPHATE PENTAHYDRATE)

24.5 mg MSP has volume 0.723 mL

in the given injection the strength of MSP is = (0.723/5) * 100 = 14.46% (v/v)